OpenStudy (anonymous):

find the surface area of a cylinder with a height of 7 inches and a radius of 3 inches. use circumference = 3.14, and round the answer to the nearest whole number

OpenStudy (anonymous):

Surface area = 9*pi + 42*pi = 51*pi (given pi = 3.14, you can calculate the rest).

Directrix (directrix):

>>circumference = 3.14 Just guessing but I think you are using the word "circumference" in place of the number "pi"

OpenStudy (anonymous):

yes pi

Directrix (directrix):

Surface area of cylinder = 2 π r h + 2 π r ² 2 π r h is the lateral (side) area and 2 π r ² is the sum of the areas of the two circular bases.

Directrix (directrix):

Surface area of cylinder = 2 π r h + 2 π r ² = 2 π (3) (7) + 2 π 3 ² = 42 π + 18 π = 60 π = 60 ( 3.14) = 188.4 approx

Directrix (directrix):

The nearest whole number would be 188 but I did not put in 3.14 for pi until the end. Just a second and I'll see if it made a difference in the whole number answer.

Directrix (directrix):

I did 42 (3.14) + 18 (3.14) and got 188.4 which is 188 to the nearest whole number.

OpenStudy (anonymous):

Oh yeah.....I forgot to add one more "area of the circle". Just added only one. Sorry.